# Model Predictive Control via Probabilistic Inference: A Tutorial and Survey

### v2

@@ 3.6.2 Convergence Analysis
Yi et al.  [yi2024covo] proved that MPPI converges for quadratic costs in the infinite-sample limit, with the convergence rate depending on the temperature parameter and the prior covariance. They further suggested that contraction properties extend to strongly convex costs with nonlinear dynamics.
Because MPPI is a zeroth-order method, its convergence is intimately tied to sample complexity. Yoon et al.  [Yoon2022-jb] and Tao et al.  [tao2022path] derived probabilistic bounds on the estimation error of the optimal distribution, showing that smaller prior covariance and lower expected trajectory costs both reduce the number of samples needed. However, convergence depends jointly on sample count, temperature, prior covariance, and cost landscape in ways that are not yet fully disentangled.
@@ 3.7 Connection to Other Paradigms
PI-MPC does not exist in isolation; it shares deep structural connections with reinforcement learning, diffusion models, and game-theoretic control. Understanding these connections clarifies which tools from neighboring fields can be imported and where PI-MPC offers unique advantages. We highlight three principal threads: the Control-as-Inference bridge to RL, the emerging link between MPPI and diffusion processes, and applications to broader problem settings.
@@
Control as Inference and Reinforcement Learning. The Control-as-Inference framework [levine2018reinforcement] provides a unified probabilistic perspective that encompasses both MPC and RL. Within this framework, RL algorithms such as Soft Actor-Critic (SAC) [haarnoja2018soft] and Maximum a Posteriori Policy Optimization (MPO) [abdolmaleki2018maximum] emerge from the same KL minimization that underlies PI-MPC; the key difference is that RL methods amortize the optimization over a learned policy, whereas PI-MPC solves it online at each time step. Model-based RL bridges the two: methods such as PETS [chua2018deep], Dreamer [hafner2019learning], and TD-MPC [hansen2022temporal, hansentd] learn dynamics and reward models while using PI-MPC planners (CEM or MPPI) for action selection. In this context, the stochastic nature of PI-MPC naturally provides the exploration variability that model-based RL requires.
@@
Diffusion Models. MPPI and score-based diffusion processes [ho2020denoising] are closely related: both iteratively refine a noisy distribution toward a target. Pan et al.  [pan2024model] and Xue et al.  [xue2024full] showed that when an optimal control problem replaces training data as the target, MPPI computations can substitute for portions of the score function, enabling sample generation at each diffusion step without training. This diffusion-style annealing captures finer modal structure than single-pass MPPI and has demonstrated strong results on high-dimensional locomotion tasks. Li et al.  [li2025unifying] further attempt to unify MPPI, RL, and diffusion models under the lens of gradient ascent on smoothed energy with respect to Gibbs measures, offering a common theoretical foundation.
Other Applications. Beyond standard single-agent control, MPPI has been extended to a range of broader settings, including multi-agent and adversarial formulations. Streichenberg et al.  [streichenberg2023multi] study MPPI in multi-agent control, and Zhang et al.  [zhang2025zsg] reformulate stochastic zero-sum games as optimal control problems that can be solved with MPPI. MPPI has also been used for data augmentation in visual navigation [inglin2026less]. Taken together, these extensions suggest that the PI-MPC framework is broadly applicable whenever sequential decision-making can be cast as trajectory optimization.
@@ 4 Conclusion
This paper has presented a tutorial and survey on probabilistic inference-based Model Predictive Control (PI-MPC). In the tutorial, we showed that the optimal control distribution admits a simple and interpretable form, i.e., the product of a Boltzmann distribution and a prior, and that practical controllers such as MPPI can be derived as variational approximations to this distribution. In the survey, we organized the growing body of PI-MPC research around the key design choices that practitioners face: how to shape the prior distribution for efficient sampling, how to handle multi-modal and constrained optimal control distributions, how to scale to high-dimensional robotic systems, and how to accelerate computation for real-time deployment.

### v3

@@ 3.6.2 Convergence Analysis
yi2024covo proved that MPPI converges for quadratic costs in the infinite-sample limit, with the convergence rate depending on the temperature parameter and the prior covariance. They further suggested that contraction properties extend to strongly convex costs with nonlinear dynamics.
Because MPPI is a zeroth-order method, its convergence is intimately tied to sample complexity. Yoon2022-jb and tao2022path derived probabilistic bounds on the estimation error of the optimal distribution, showing that smaller prior covariance and lower expected trajectory costs both reduce the number of samples needed. However, convergence depends jointly on sample count, temperature, prior covariance, and cost landscape in ways that are not yet fully disentangled.
@@ 3.7 Connection to Other Paradigms
PI-MPC does not exist in isolation; it shares deep structural connections with RL, diffusion models, and data-driven learning. Understanding these connections clarifies which tools from neighboring fields can be imported and where PI-MPC offers unique advantages. We highlight three principal threads: the Control-as-Inference bridge to RL, the emerging link between MPPI and diffusion processes, and the use of PI-MPC as an engine for data generation and policy learning.
Control as Inference and Reinforcement Learning.
@@
The Control-as-Inference framework (levine2018reinforcement) provides a unified probabilistic perspective that encompasses both PI-MPC and RL. Within this framework, RL algorithms such as Soft Actor-Critic (SAC) (haarnoja2018soft) and Maximum a Posteriori Policy Optimization (MPO) (abdolmaleki2018maximum) emerge from the same KL-regularized inference that underlies PI-MPC; the key difference is that RL methods amortize the optimization over a learned policy, whereas PI-MPC solves it online at each time step. Model-based RL bridges the two: methods such as PETS (chua2018deep), Dreamer (hafner2019learning), and TD-MPC (hansen2022temporal; hansentd) learn dynamics and reward models while using PI-MPC planners (CEM or MPPI) for action selection. In this context, the stochastic nature of PI-MPC naturally provides the exploration variability that model-based RL requires. Empirically, wangbootstrapped report that, on more challenging tasks, incorporating MPPI into a model-based RL pipeline leads to more robust value-function learning and improved control performance over a model-free SAC-style policy alone.
Diffusion Models.
@@
MPPI and score-based diffusion processes (ho2020denoising) are closely related: both iteratively refine a noisy distribution toward a target. pan2024model and xue2024full showed that when an optimal control problem replaces training data as the target, MPPI computations can substitute for portions of the score function, enabling sample generation at each diffusion step without training. This diffusion-style annealing captures finer modal structure than single-pass MPPI and has demonstrated strong results on high-dimensional locomotion tasks. li2025unifying further attempt to unify MPPI, RL, and diffusion models under the lens of gradient ascent on smoothed energy with respect to Gibbs measures, offering a common theoretical foundation.
Data Generation.
The stochastic nature of PI-MPC also makes it a natural fit for data augmentation and sample-efficient learning. In visual navigation, inglin2026less use MPPI rollouts to augment training data for imitation learning-based navigation. PI-MPC can further be coupled with learned dynamics models to generate large amounts of synthetic experience: morgan2021model run MPPI on a learned world model (ha2018world) to produce model-generated trajectories that are mixed with real-environment data for off-policy RL, while wangcoplanner design a dual strategy in which the world model is trained by sampling transitions that reduce epistemic uncertainty, whereas the MPPI cost function is modified at deployment to actively steer the agent into uncertain regions for more informative data collection. While the above approaches use MPPI to produce training data, it can also guide policy learning directly: wangbootstrapped distill the Gaussian action distribution of an MPPI planner into a model-free policy, achieving high training efficiency by bootstrapping model-free RL with model-based learning and control.
@@ 4 Conclusion
This paper has presented a unified tutorial and survey on probabilistic inference-based Model Predictive Control (PI-MPC). In the tutorial, we showed that the optimal control distribution admits a simple and interpretable form, i.e., the product of a Boltzmann distribution and a prior, and that practical controllers such as MPPI can be derived as variational approximations to this distribution. In the survey, we organized the growing body of PI-MPC research around the key design choices that practitioners face: how to shape the prior distribution for efficient sampling, how to handle multimodal and constrained optimal control distributions, how to scale to high-dimensional robotic systems and real-time deployment, and how PI-MPC connects to related theoretical and learning-based paradigms.
Future directions.
Given the recent success of learning-based control, e.g., large-scale reinforcement learning, PI-MPC can offer additional value as a complementary component rather than a competing alternative. Because PI-MPC can be connected to reinforcement learning and diffusion-based generative processes through a unified variational inference framework, as discussed in Section 3.7, it may provide a useful lens for analyzing these methods and for developing clearer theoretical interpretations and guarantees. Moreover, purely learning-based controllers often require extensive training and can be less flexible in incorporating safety constraints or structured environment knowledge. Hybrid approaches that integrate PI-MPC with learning-based methods therefore have the potential to combine the strengths of both paradigms: fast online adaptation and constraint-aware planning from PI-MPC, together with representation learning and generalization from data-driven models.